OpenStudy (anonymous):

Find the exponential function f(x)=a^x whose graph goes through the point (3,125). a = ??

OpenStudy (anonymous):

a=5

OpenStudy (anonymous):

125=a^3 5^3=a^3 a=3

OpenStudy (anonymous):

sorry 5